HIGH RESOLUTION INFRARED SPECTROSCOPY OF THE FLUOROMETHYL RADICAL

\underline{E.S. WHITNEY}, F. DONG, D.J. NESBITT, JILA and National Institute of Standards and Technology, University of Colorado, Boulder, CO 80309.

The CH_{2}F radical is an intermediate in the decomposition of various atmospheric species and is also important in free radical addition reactions. Previous work on the CH_{2}F radical has utilized microwave spectroscopy to determine rotational transitions within the ground vibrational state (the symmetric C-H stretch) ^{a}. Infrared spectroscopic methods have also been used to probe the v3 band (C-F stretch) origin and rotational constants ^{b}. We have used high resolution (0.0005 cm^{−1}) infrared spectroscopy to investigate rovibrational transitions and assign band origins for both the symmetric and asymmetric stretch regions of the CH_{2}F radical. Measurements were taken using direct absorption (sensitivity = 1 x 10^{−7} per root Hz) under slit-jet cooled conditions (20 K). Assignment of the spectrum (up to J=7) has yielded refined lower and upper state rotational and centrifugal distortion constants for this near-prolate top. The ratio of absorption intensities for the symmetric and asymmetric bands provides an interesting comparison to that of the CH_{2}Cl radical, where no asymmetric band was detected. Interpretation of hyperfine structure will also be discussed.

aY. Endo, C. Yamada, S. Saito, and E. Hirota, J. Chem. Phys. 79(4), 1983.

bC. Yamada and E. Hirota, J. Molec. Spec. 116, 1986.